solve algebraically for the value of x where these two lines cross:3y=21-6x and 3y=24-9x (show work)
Answer by rapaljer(4671)

Since you have "3y = ___" in both equations, and since the instructions say to solve algebraically for x, just set them equal to each other, and solve for x:
3y = 3y
21-6x = 24- 9x

Add +9x to each side to get the x terms on the left side:
21 -6x + 9x = 24 - 9x + 9x
21+3x = 24

Next, subtract 21 from each side:
21-21 + 3x = 24 - 21
3x = 3

Divide both sides by 3:
x = 1.

Check your answer:
21 - 6x = 24-9x
21 -6(1) = 24 - 9(1)
15 = 15
It checks!

R^2 at SCC